OpenStudy (anonymous):

Another u-substitution integration.. : sin^6(5x)cos(5x)dx

OpenStudy (anonymous):

u = 5x, du = 5dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin(u)^6 \cos(u) \frac{ du }{ 5 }\]

OpenStudy (anonymous):

the answer is \[\frac{ 1 }{ 35 }\sin^7 (5x) + C\]

OpenStudy (anonymous):

but im not sure how to eliminate the \[\cos(u)\] in the integral

ganeshie8 (ganeshie8):

try : \(u = \sin(5x)\)

OpenStudy (anonymous):

alright, let's see

OpenStudy (anonymous):

would the derivative be -cos5x?

ganeshie8 (ganeshie8):

\(\frac{d}{dx}\sin x = \cos x\)

OpenStudy (anonymous):

\[\int\limits_{}^{}(u)^6 (du)\]

ganeshie8 (ganeshie8):

dont forget the chain rule

ganeshie8 (ganeshie8):

u dont have just sinx, u are substituting sin(5x)

OpenStudy (anonymous):

5cos(5x)

ganeshie8 (ganeshie8):

\(u = \sin 5x\) \( du = 5\cos (5x) dx \)

OpenStudy (anonymous):

\[\int\limits_{}^{}(u)^6 5du \]

ganeshie8 (ganeshie8):

not quite, lets do it step by step

OpenStudy (anonymous):

alright, sorry and thanks!!

ganeshie8 (ganeshie8):

\(u = \sin 5x\) \( du = 5\cos (5x) dx \) \(\implies \) \( \dfrac{du}5{} = \cos (5x) dx \)

ganeshie8 (ganeshie8):

\[\int \sin^6(5x) \cos(5x)dx \] becomes : \[\int u^6 \frac{du}{5} \]

OpenStudy (anonymous):

oh, okay I understand

OpenStudy (anonymous):

is it always safe to assume that: du = 1?

OpenStudy (anonymous):

i got the answer btw, just wanted to know ^

ganeshie8 (ganeshie8):

hmm what do u mean assume `du = 1` ?

OpenStudy (anonymous):

hm, \[\int\limits_{}^{}(u)^6 (\frac{ du }{ 5 })\] becomes

OpenStudy (anonymous):

\[(\frac{ 1 }{ 7 }u^7) (\frac{ 1 }{ 5 }) + C \]

ganeshie8 (ganeshie8):

^thats right, i think i got what you're asking

ganeshie8 (ganeshie8):

\(\color{red}{\int} \) is meaningless without the differential \(\color{red}{dx}\) or \(\color{red}{du}\)

ganeshie8 (ganeshie8):

below "entire thing" is called the integral operator : \(\color{red}{\int~dx} \)

ganeshie8 (ganeshie8):

you can think of \(\color{red}{\int}\) as the "starting parenthesis", and the \(\color{red}{dx}\) or \(\color{red}{du}\) as the "ending parenthesis"

OpenStudy (anonymous):

ah, that makes a lot of sense. Thanks so much! i would give double medal if I could!

ganeshie8 (ganeshie8):

And, the stuff that lies in between is called the `integrand`, which you will be integrating : \(\color{red}{\int}\) `integrand`\(\color{red}{dx}\)

ganeshie8 (ganeshie8):

np... u wlc :)